Evaluate the functions for the values of x given as 1, 2, 4, 8, and 16. Describe the differences in the rate at which each function changes with increasing values of x.
f(x) = 2x – 5 Thanks!

Answer by RAY100(1637)

f(x) = 2x-5
.
x,,,,,,f(x),,,,,,,,,,,,,,,,,,,,Difference / Unit ,,,,,,,,,,,,,,,,,,,,,,rate
.
1,,,,,,2(1)-5 = -3
.,,,,,,,,,,,,,,,,,,,,,,,,,,,,,,,,,,,,,,,,,,,2/1,,,,,,,,,,,,,,,,,,,,,,,,,,,,,,,2
2,,,,,,,2(2) -5 = -1
.,,,,,,,,,,,,,,,,,,,,,,,,,,,,,,,,,,,,,,,,,,,,4/2,,,,,,,,,,,,,,,,,,,,,,,,,,,,,,2
4,,,,,,,,2(4)-5 = 3
.,,,,,,,,,,,,,,,,,,,,,,,,,,,,,,,,,,,,,,,,,,,,8/4,,,,,,,,,,,,,,,,,,,,,,,,,,,,,,2
8,,,,,,,,2(8)-5 = 11
.,,,,,,,,,,,,,,,,,,,,,,,,,,,,,,,,,,,,,,,,,,,,,16/8,,,,,,,,,,,,,,,,,,,,,,,,,,,2
16,,,,,,2(16) -5 = 27
.
Rate of change is 2 throughout
.
Arithmetic form is a(n) = a(1) +(n-1) d
.
comparing to f(x) = 2x-5 has,,, a(0) = -5, d=2